1. You are studying a color polymorphism in a population of soapberry bugs. A single gene controls color variation; individuals with at least 1 copy of the R allele at this locus are red while individuals of the rr genotype are white. You know that this population satisfies the conditions for Hardy-Weinberg Equilibrium, and that 64% of the population shows the red phenotype. What is the frequency of the R allele? Show your work.

2. Male Trinidadian guppies use bright red patches of color to attract mates. Your colleague has recently sequenced 200 individual guppies at the opsin gene that controls sensitivity to red wavelengths of light. He has found a new variant of this gene that increases sensitivity to red in his samples. Unfortunately, a computer malfunction has scrambled his data and he no longer knows which alleles came from heterozygotes or homozygotes. Based on his records, 40 of the 400 alleles sequenced are the new variant. If this gene is in Hardy-Weinberg Equilibrium in this population, how many individuals from his sample of 200 should be homozygous for the new allele?
Answer #1

percentage of red soapberry bugs= 64%

percentage of white soapberry bugs= 100-64=36%

the genotype of white soapberry bugs is rr.

genotypic frequency of rr= 36/100=0.36

allelic frequency of r= genotypic frequency of rr

=0.36

= 0.6

given population is in Hardy-Weinberg equilibrium

so allelic frequency of R + allelic frequency of r=1

allelic frequency of R= 1-allelic frequency of r

= 1- 0.6

= 0.4

frequency of the R allele =0.4